### v2

@@ 1 Introduction
Hertog et al. Hertog:2018kbz examined the stability of the classic Euclidean Giddings–Strominger axion-sourced wormhole Giddings:1987cg , which prompted much of the classic work on wormhole-related apparent paradoxes. To a sigh of relief, they appeared to show that the solutions were simply unstable, at least in 44D asymptotically flat spacetime or AdS4 with a large AdS radius relative to the wormhole mouth radius. However, subsequent analyses Loges:2022nuw revealed that in the dual description of the axion as a $2$-form field strength, the wormholes are actually stable in flat space. The source of the discrepancy, as fully elucidated later in Hertog:2024nys , is a difference in boundary conditions: authors of Hertog:2018kbz implicitly chose a Robin boundary condition. For boundary conditions which fix the axion charge and boundary metric, it is now clear from Loges:2022nuw ; Hertog:2024nys that the axion wormholes are stable, in spite of other boundary conditions leading to instability.
@@
In 33D, the axion field responsible for Giddings–Strominger wormholes is equivalent to a $U(1)$ gauge field, the “dual photon” TongGaugeTheory . This provides an especially tractable situation with well-studied holographic applications Marolf_2006 , enabling detailed analysis of wormhole contributions and associated paradoxes. That is precisely the aim of this paper, where we extend previous results Hertog:2018kbz ; Loges:2022nuw ; Hertog:2024nys ; Barcelo:1995gz to AdS3 backgrounds, for arbitrary ratio of the wormhole mouth radius to the AdS radius, and to torus and compact hyperbolic topologies, placing wormholes in a context whose implications may be more theoretically tractable.
@@ 3.3 Zero modes
It is by no means obvious that this is postive definite, since there is a large negative contribution from deep within the wormhole throat $T\to 0$. However, we have not yet implemented the boundary conditions. As explained in Marolf_2006 , and which will important for interpreting the solutions, the boundary conditions for $D=3$ are to fix the field strength at the boundary, not (as may be expected) the surface potential. The condition may be written covariently as
$\lim_{T\to\pm\infty}\sqrt{h}n^{\mu}F_{\mu i}=0,$ (50)
where $n^{\mu}$ is a unit vector orthogonal to the boundary. Implementing this yields
@@ 3.4 Boundary Conditions and UV completions
It is important to note that the boundary conditions 50, while different than the “usual” for a $U(1)$ gauge field, are identical to those implemented in Loges:2022nuw ; Hertog:2024nys . The idea, reviewed nicely in Loges:2022nuw ; Hertog:2024nys , is that Dirichlet boundary conditions on the field strength correspond to the (standard) Neumann boundary conditions for the dual axion. Indeed, this fact was referenced in Marolf_2006 as one way of understanding the meaning of the boundary condition for a gauge field in $D=3$. So our conditions are the same as Loges:2022nuw ; Hertog:2024nys , so our solutions are the $AdS_{3}$ analogues of theirs, and we likewise find them to be stable.
@@
Alternatively, one could attempt to take a top-down perspective on interpreting our stability results. Indeed, the setting of Arkani-Hamed:2007cpn which connected axion wormholes to CFT cluster-decomposition considered the embedding of axion wormholes in a toriodal compactification of Type IIB String Theory to $AdS_{3}$, and axions of some sort abound in String Theory compactifications generally Svrcek:2006yi . However, there are many barriers to UV embedding of our model. For example, one expects there to be dilaton fields coupled to any axions, and in some cases Astesiano this has precluded embeddings while in others solutions remained and stability was retained Hertog:2024nys . One also typically expects Chern-Simons terms for a gauge field (in the axion frame, this is related to the Peccei–Quinn theory cosine potential) or charged fields which could trigger the Polyakov confining mechanism Polyakov:1975rs . It would be interesting to know if these obstacles can be overcome, or alternatively how each effect the existence and stability of solutions in generality, since this may offer answers to any associated wormhole paradoxes. Even more provocatively, the tensionless limit of string theory was investigated in $AdS_{3}$ and it was argued that including wormhole backgrounds was redundant and they should not be included as distinct saddles at all Eberhardt:2020bgq ; Eberhardt:2021jvj .
@@ 5 Discussion
In this paper we have found flat, positively curved, and negatively cured analogues of the Giddings–Strominger “axion” Euclidean wormholes in the context of asymptotically AdS3 Einstein gravity with a $U(1)$ field, found that they are stable, and in some cases computed their classical action. We have done so by exploiting some convenient properties of the ADM formalism in the context of $D=3$ gravity. We have also found a convenient set of coordinates that cover the entirety of the two-sided solutions and which make manifest their nonsingularity.
This builds on the recent work of Hertog:2018kbz ; Loges:2022nuw ; Hertog:2024nys showing yet another class of stable wormhole contributions to the gravitational path integral that heighten the associated paradoxes. It therefore seems increasingly likely that any hope that the problematic wormholes are simply unstable is unrealistic, however it is important to assess this in the context of a UV complete theory such as String Theory. Including dilaton fields, or better yet considering a construction such as those considered in Arkani-Hamed:2007cpn and reexamining stability would provide stronger evidence. Some nice recent work along these lines has made progress Andriolo:2022rxc

### v3

@@ 1 Introduction
Hertog et al. Hertog:2018kbz examined the stability of the classic Euclidean Giddings–Strominger axion-sourced wormhole Giddings:1987cg , which prompted much of the classic work on wormhole-related apparent paradoxes. To a sigh of relief, they appeared to show that the solutions were simply unstable, at least in 44D asymptotically flat spacetime or AdS4 with a large AdS radius relative to the wormhole mouth radius. However, subsequent analyses Loges:2022nuw revealed that in the dual description of the axion as a $2$-form field strength [^1] While this manuscript was undergoing revision, the very relevant witten2026dualityaxionwormholes appeared, offering a systematic assessment of this duality of descriptions and associated subtleties., the wormholes are actually stable in flat space. The source of the discrepancy, as fully elucidated later in Hertog:2024nys , is a difference in boundary conditions: authors of Hertog:2018kbz implicitly chose a Robin boundary condition. For boundary conditions which fix the axion charge and boundary metric, it is now clear from Loges:2022nuw ; Hertog:2024nys that the axion wormholes are stable, in spite of other boundary conditions leading to instability.
@@
In 33D, the axion field responsible for Giddings–Strominger wormholes is equivalent to a $U(1)$ gauge field, the “dual photon” TongGaugeTheory .  [^2] Although, importantly, the boundary conditions must be chosen differently in $D=3$, see section 3.4. This provides an especially tractable situation with well-studied holographic applications Marolf_2006 , enabling detailed analysis of wormhole contributions and associated paradoxes. That is precisely the aim of this paper, where we extend previous results Hertog:2018kbz ; Loges:2022nuw ; Hertog:2024nys ; Barcelo:1995gz to AdS3 backgrounds, for arbitrary ratio of the wormhole mouth radius to the AdS radius, and to torus and compact hyperbolic topologies, placing wormholes in a context whose implications may be more theoretically tractable.
@@ 3.3 Zero modes
It is by no means obvious that this is postive definite, since there is a large negative contribution from deep within the wormhole throat $T\to 0$. However, we have not yet implemented the boundary conditions. As explained in Marolf_2006 , and which is important for interpreting the solutions, the boundary conditions for $D=3$ are to fix the field strength at the boundary, not (as may be expected) the surface potential. The condition may be written covariently as
$\lim_{T\to\pm\infty}\sqrt{h}n^{\mu}F_{\mu i}=constant,$ (50)
where $n^{\mu}$ is a unit vector orthogonal to the boundary. For the fluctuations to preserve the boundary condition, the constant must be =0=0. Implementing this yields
@@ 3.4 Boundary Conditions and UV completions
It is important to note that the boundary conditions 50, while the correct and only possible choice for $D=3$, are different from both the “usual” for a $U(1)$ gauge field and from those implemented in Loges:2022nuw ; Hertog:2024nys .
The boundary conditions in this work, fixing the normal component of the field strength, correspond to fixing the $U(1)$ electric charge (by Gauss’s law). As remarked on in Marolf_2006 , this is different than the more common choice of fixing the boundary potential, and this is significant because only in the latter case is the bulk gauge field dual to a CFT conserved charge. This matters for the proper CFT interpretation of these wormholes, e.g. which/which kind of operators may be expected to potentially exhibit related effects.
In Loges:2022nuw ; Hertog:2024nys the tangential component of the gauge field is fixed. This corresponds to fixing the value of the axion charge, which in $D=3$ could be referred to as the dual photon charge. As explained in Marolf_2006 and more recently in Held:2026huj , this is not allowed in $D=3$ and a simple way to see this is that the CFT operator dual to a scalar of mass $m$ in dimension $d$ is given by $\Delta_{\pm}=\frac{d}{2}\pm\sqrt{\frac{d^{2}}{4}+m^{2}L^{2}}$ where +/- maps to Dirichlet/Neumann. For Neumann for $m=0$ (the usual axion case) the CFT would have an operator of dimension $\Delta_{-}=0$ which violates the unitarity bound. So this is not allowed.
Nevertheless, interestingly, for the case of the sphere which is the direct analogue of Loges:2022nuw ; Hertog:2024nys (and the hyperbolic quotients) the stability did not depend in any way on the boundary conditions, meaning that even if one were to consider an (ultimately unphysical) toy model which is the direct analogue of their solutions, it would likewise exhibit the same stability.
@@
Alternatively, one could attempt to take a top-down perspective on interpreting our stability results. Indeed, the setting of Arkani-Hamed:2007cpn which connected axion wormholes to CFT cluster-decomposition, considered the embedding of axion wormholes in a toriodal compactification of Type IIB String Theory to $AdS_{3}$, and axions of some sort abound in String Theory compactifications generally Svrcek:2006yi . However, there are many barriers to UV embedding of our model. For example, one expects there to be dilaton fields coupled to any axions, and in some cases Astesiano this has precluded embeddings while in others solutions remained and stability was retained Hertog:2024nys . One also typically expects Chern-Simons terms for a gauge field (in the axion frame, this is related to the Peccei–Quinn theory cosine potential) or charged fields which could trigger the Polyakov confining mechanism Polyakov:1975rs . It would be interesting to know if these obstacles can be overcome, or alternatively how each effect the existence and stability of solutions in generality, since this may offer answers to any associated wormhole paradoxes. Even more provocatively, the tensionless limit of string theory was investigated in $AdS_{3}$ and it was argued that including wormhole backgrounds was redundant and they should not be included as distinct saddles at all Eberhardt:2020bgq ; Eberhardt:2021jvj .
@@ 5 Discussion
In this paper we have found flat, positively curved, and negatively curved analogues of the Giddings–Strominger “axion” Euclidean wormholes in the context of asymptotically AdS3 Einstein gravity with a $U(1)$ field, found that they are stable, and in some cases computed their classical action. We have done so by exploiting some convenient properties of the ADM formalism in the context of $D=3$ gravity. We have also found a convenient set of coordinates that cover the entirety of the two-sided solutions and which make manifest their nonsingularity.
This builds on the recent work of Hertog:2018kbz ; Loges:2022nuw ; Hertog:2024nys showing yet another class of stable wormhole contributions to the gravitational path integral that heighten the associated paradoxes (even as the allowed boundary conditions in $D=3$ mean our solutions differ in some important respects). It therefore seems increasingly likely that any hope that the problematic wormholes are simply unstable is unrealistic, however it is important to assess this in the context of a UV complete theory such as String Theory. Including dilaton fields, or better yet considering a construction such as those considered in Arkani-Hamed:2007cpn and reexamining stability would provide stronger evidence. Some nice recent work along these lines has made progress Andriolo:2022rxc
@@
While this manuscript was undergoing revision, the very interesting, aforementioned, Held:2026huj also investigated $AdS_{3}$ axion wormholes. It would be very interesting to do a full comparison of results.